# A Style-Based Metric for Quantifying the Synthetic-to-Real Gap in Autonomous Driving Image Datasets

### v1

@@ Abstract
Ensuring the reliability of autonomous driving perception systems requires extensive environment-based testing, yet real-world execution is often impractical. Synthetic datasets have therefore emerged as a promising alternative, offering advantages such as cost-effectiveness, bias free labeling, and controllable scenarios. However, the domain gap between synthetic and real-world datasets remains a critical bottleneck for the generalization of AI-based autonomous driving models. Quantifying this synthetic-to-real gap is thus essential for evaluating dataset utility and guiding the design of more effective training pipelines. In this paper, we establish a systematic framework for quantifying the synthetic-to-real gap in autonomous driving systems, and propose Style Embedding Distribution Discrepancy (SEDD) as a novel evaluation metric. Our framework combines Gram matrix-based style extraction with metric learning optimized for intra-class compactness and inter-class separation to extract style embeddings. Furthermore, we establish a benchmark using publicly available datasets. Experiments are conducted on a variety of datasets and sim-to-real methods, and the results show that our method is capable of quantifying the synthetic-to-real gap. This work provides a standardized quality control tool that enables systematic diagnosis and targeted enhancement of synthetic datasets, advancing future development of data-driven autonomous driving systems.
@@ I Introduction
To address these issues, we introduce a novel unified synthetic-to-real gap quantification model with a style-based metric, Style Embedding Distribution Discrepancy (SEDD), which aims to offer a general solution for quantifying the synthetic-to-real gap. The main contributions of this work are as follows:
  • We innovatively model the discrepancies between synthetic and real data as differences in style and introduce a data fidelity metric Style Embedding Distribution Discrepancy (SEDD).
@@
  • A large number of experiments have been conducted on real datasets, synthetic datasets to demonstrate the effectiveness of our proposed quantification model.
  • We apply the proposed model to sim-to-real approaches in autonomous driving, providing quantitative metrics and visualization tools for their evaluation.
@@ II-C Synthetic-to-Real Gap Quantification for Autonomous Driving
Figure 2: The overall framework of our proposed synthetic-to-real gap quantification model. The model takes as input real images, synthetic images under various weather conditions. The input images are sequentially processed through three key modules: feature extractor, style extractor, and metric learning module. During training, the model parameters are optimized using a combination of Center Loss and NTXent Loss. In the evaluation phase, the learned feature embeddings are post-processed to compute the final SEDD metric.
@@ III Methodology
The overall framework of our proposed model is illustrated in Fig. 2. The model takes as input real images, synthetic images under various weather conditions. During the training phase, input data is processed through a feature extractor to obtain feature maps, followed by a style extractor that generates feature embeddings. These embeddings are then processed using metric learning, incorporating Center Loss and NTXent Loss. The combined loss is propagated backward to update the parameters of the feature extractor and style extractor. In the evaluation phase, the learned feature embeddings undergo post-processing to compute fidelity metrics, providing an objective quantification of the synthetic-to-real gap.
@@ III-E Post-Processing
Figure 4: Visualization of results on validation set and for sim-to‑real methods. The result visually demonstrates that our quantification model can distinguish Virtual KITTI dataset from Virtual KITTI 2 dataset. And the samples after sim-to-real are closer to the real ones.
@@ IV-B1 Quantification Results on Validation Set
The results of our trained model on the validation set are shown in Table I. This result provides a reference value for subsequent quantification. Notably, during the training process, no label prior information indicating that Virtual KITTI 2 is superior was provided. Nevertheless, our model autonomously determines that the style of Virtual KITTI 2 is closer to that of the real dataset, resulting in a smaller synthetic-to-real gap. Fig. 4a presents the visualization results. In the figure, the sample points are clustered by color rather than by shape, which indicates that our method achieves the disentanglement of content and style.
@@ IV-C Ablation Study
We evaluated the effect of the loss function using the internal variance of the real dataset. As shown in Table V, the experimental results indicate that using a single loss function leads to an increase in the intra-class variance of the feature distributions, whereas the joint loss effectively constrains the compactness of intra-class samples. The ablation experiments further demonstrate that solely relying on $\mathcal{L}_{\text{NTXent}}$ may cause the model to overlook intra-class consistency, highlighting the need for an explicit intra-class constraint. In contrast, optimizing $\mathcal{L}_{C}$ in isolation tends to converge to a local optimum, while the inter-class information provided by the contrastive loss helps mitigate it.
@@ V Conclusion
In this work, we propose a novel unified synthetic-to-real gap quantification model with a style-based metric, Style Embedding Distribution Discrepancy (SEDD). Using a ResNet-based feature extractor, our approach derives style embeddings through Gram matrices and optimizes them with metric learning, enabling similarity quantification between synthetic and real datasets. Experiments validate the effectiveness of the proposed metric. This work transforms the synthetic-to-real gap into quantifiable objective indicators, eliminating reliance on subjective experience. It provides a standardized measurement tool for the quality control of synthetic datasets in autonomous driving, facilitating more targeted improvements. Future work could extend this framework to generative model-based synthetic datasets, such as world model related work.

### v2

# A Style-Based Profiling Framework for Quantifying the Synthetic-to-Real Gap in Autonomous Driving Datasets

@@ Abstract
Ensuring the reliability of autonomous driving perception systems requires extensive environment-based testing, yet real-world execution is often impractical. Synthetic datasets have therefore emerged as a promising alternative, offering advantages such as cost-effectiveness, bias free labeling, and controllable scenarios. However, the domain gap between synthetic and real-world datasets remains a major obstacle to model generalization. To address this challenge from a data-centric perspective, this paper introduces a profile extraction and discovery framework for characterizing the style profiles underlying both synthetic and real image datasets. We propose Style Embedding Distribution Discrepancy (SEDD) as a novel evaluation metric. Our framework combines Gram matrix-based style extraction with metric learning optimized for intra-class compactness and inter-class separation to extract style embeddings. Furthermore, we establish a benchmark using publicly available datasets. Experiments are conducted on a variety of datasets and sim-to-real methods, and the results show that our method is capable of quantifying the synthetic-to-real gap. This work provides a standardized profiling-based quality control paradigm that enables systematic diagnosis and targeted enhancement of synthetic datasets, advancing future development of data-driven autonomous driving systems.
@@ I Introduction
To address these issues, we introduce a novel style-based profiling framework with a metric, Style Embedding Distribution Discrepancy (SEDD). The main contributions of this work are as follows:
  • We innovatively model the discrepancies between synthetic and real data as differences in extracted style profiles and propose a data fidelity metric Style Embedding Distribution Discrepancy (SEDD).
@@
  • A large number of experiments have been conducted on real datasets, synthetic datasets to demonstrate the effectiveness of our proposed profiling framework.
  • We apply the proposed profiling framework to sim-to-real approaches in autonomous driving, providing quantitative metrics and visualization tools for their evaluation.
@@ II-C Synthetic-to-Real Gap Quantification for Autonomous Driving
Figure 2: The overall profiling framework. The model takes as input real images, synthetic images under various weather conditions. The input images are sequentially processed through three key modules: feature extractor, style extractor, and metric learning module. During training, the model parameters are optimized using a combination of Center Loss and NTXent Loss. In the evaluation phase, the learned feature embeddings are post-processed to compute the final SEDD metric.
@@ III Methodology
The overall profiling framework is illustrated in Fig. 2. The framework takes as input real images, synthetic images under various weather conditions. During the training phase, input data is processed through a feature extractor to obtain feature maps, followed by a style extractor that generates feature embeddings. These embeddings are then processed using metric learning, incorporating Center Loss and NTXent Loss. The combined loss is propagated backward to update the parameters of the feature extractor and style extractor. In the evaluation phase, the learned feature embeddings undergo post-processing to compute fidelity metrics, providing an objective quantification of the synthetic-to-real gap.
@@ III-E Post-Processing
Figure 4: Visualization of results on validation set and for sim-to‑real methods. The result visually demonstrates that our profiling framework can distinguish Virtual KITTI dataset from Virtual KITTI 2 dataset. And the samples after sim-to-real are closer to the real ones.
@@ IV-B1 Quantification Results on Validation Set
The results of our trained profiling framework on the validation set are shown in Table I. This result provides a reference value for subsequent quantification. Notably, during the training process, no label prior information indicating that Virtual KITTI 2 is superior was provided. Nevertheless, our profiling framework autonomously determines that the style of Virtual KITTI 2 is closer to that of the real dataset, resulting in a smaller synthetic-to-real gap. Fig. 4a presents the visualization results. In the figure, the sample points are clustered by color rather than by shape, which indicates that our method achieves the disentanglement of content and style.
@@ IV-C Ablation Study
We evaluated the effect of the loss function using the internal variance of the real dataset. As shown in Table V, the experimental results indicate that using a single loss function leads to an increase in the intra-class variance of the feature distributions, whereas the joint loss effectively constrains the compactness of intra-class samples. The ablation experiments further demonstrate that solely relying on $\mathcal{L}_{\text{NTXent}}$ may cause the framework to overlook intra-class consistency, highlighting the need for an explicit intra-class constraint. In contrast, optimizing $\mathcal{L}_{C}$ in isolation tends to converge to a local optimum, while the inter-class information provided by the contrastive loss helps mitigate it.
@@ V Conclusion
In this work, we propose a novel style-based profiling framework with a metric, Style Embedding Distribution Discrepancy (SEDD). Using a ResNet-based feature extractor, our approach derives style embeddings through Gram matrices and optimizes them with metric learning, enabling similarity quantification between synthetic and real datasets. Experiments validate the effectiveness of the proposed metric. This work transforms the synthetic-to-real gap into quantifiable objective indicators, eliminating reliance on subjective experience. It provides a standardized profiling-based tool for the quality control of synthetic datasets in autonomous driving, facilitating more targeted improvements. Future work could extend this framework to generative model-based synthetic datasets, such as world model related work.